OpenStudy (anonymous):

which of the following are solutions to the equation sinx cos(pi/7)-sin(pi/7)cosx=sqrt(2)/2 check all that apply A. 7pi/4+pi/4+2npi B. 5pi/4+pi/7+2npi C. pi/4+pi/7+2npi D. 3pi/4+pi/7+2npi

OpenStudy (anonymous):

I can not figure this out

OpenStudy (anonymous):

sinx cos(pi/7) - sin(pi/7) cosx = sqrt(2)/2 sin(x - pi/7) = sqrt(2)/2 x - pi/7 = pi/4 or x-pi/7 = 3pi/4 x = (11pi/28) + k(2pi) or x = (25pi/28) + k(2pi), where k is an integer So actually both C and D are the the solutions

OpenStudy (anonymous):

thats what i thought but then i doubted myself. thanks

OpenStudy (anonymous):

your welcome!